If Mark is 17, and Jimmy is twice Mark's age, then how old is Jimmy when Mark is 20
Found 2 solutions by ewatrrr, TimothyLamb:
Answer by ewatrrr(24785)

 
Hi,
If Mark is 17, and Jimmy is twice Mark's age, then Jimmy is 34.
In 3 years, Jimmy will be 37.

Answer by TimothyLamb(4379)

m = 17
j = 34
---
j - m = 17
j = 17 + m
---
when mark is 20:
j = 17 + 20
---
j = 37
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php